Calculate unemployment rate if adult population of the nation is 800 people of which 450 people are working and 30 are looking for working.
Answer #1

Answer:-

unemployment rate =(unemployed/labor force )*100
labor force =employed +unemployed

labor force =450+30

Labour Force= 480 People

Employed=450 People

Unemployed=Total Population -Labour force

=800-480=320

Unemployment rate =(unemployed/Total Population )*100

=(320/800)*100

=0.40*100=40%

Unemployment rate =40%